OpenStudy (anonymous):

What is the equation of direct variation, given that y varies directly with x. y is 10 when x is 2

OpenStudy (anonymous):

y =5x

OpenStudy (anonymous):

This helped me. If y varies directly as x, and y = 8 when x = 12, find k and write an equation that expresses this variation. Plan of Attack: Plug the given values into the equation y = kx. Solve for k. Then replace k with its value in the equation y = kx. Step-by-Step: Start with our standard equation: y = kx Insert our known values: 8 = k*12 Divide both sides by 12 to find k: 8/12 = k 2/3 = k Next: Go back to y = kx and replace k with 2/3. Result: y = (2/3)x

OpenStudy (anonymous):

thanks!(: